Find the following derivatives: ON Js, V.X1, X2....,XC-1 ON T,P.x1,x2,...,XC-1
Answer #1

Ant given that - ,V, r, I 2 - YC-1 and in sy to sch * Ist law of th since we know That By using - * ist law of TheymodyanuiesGibs Eneway of the system - C4 20 = du+ pdv Tds = dot pdv ? CH = H-TS dy= dH-Tds - sdt DH = Tds + valp so dey= Tds + vdp - Td